Two balls, A and B, are launched horizontally at the same time from the same height.  Ball A is launched with an initial speed of 12.6 m s-1 and lands 23.2 m away.  Ball B lands 56.4 m away.

  1. How high were the balls launched from?
  2. How fast was ball B launched at?
1
Expert's answer
2021-10-31T18:11:39-0400

The height is


"h=gt^2/2=g/2·(R_A/v_A)^2=16.6\text{ m}."

The speed of the second ball is


"v_B=R_B\sqrt{\frac{g}{2h}}=30.6\text{ m/s}."